Perform the following calculations (24 pts.) a. 20.0 g H2O (18.0 g/mol) to moles




b. 1.00 mol C6H6 (78.1 g/mol) to grams



c. 5.00 x 1022 molecules O2 (32.0 g/mol) to grams




d. 0.333 g C6H12O6 (180.2 g/mol) to moles



e. 12.2 g CCl4 (153.8 g/mol) to atoms of Cl



f. 4.20 g HNO3 (63.0 g/mol) to molecules HNO3
Answer #1

a)

Mole = mass /molar mass

moles of H2O = 20.0g/18.0g

moles of H2O = 1.111mol

b)

Mass = mole × molar mass

mass of C6H6 = 1.00mol × 78.1gmol

mass of C6H6 = 78.1g

c)

1mole = 6.022×1023molecules

moles of O2 = (1mol/6.022×1023molecules) ×5.00 ×1022 molecules = 0.083029mol

mass of O2 = 0.083029mol × 32.0g/mol

mass of O2 = 2.657g

d)

Moles of C6H12O6 = 0.333g / 180.2g/mol

moles of C6H12O6 = 0.001848mol

e)

moles of CCl4 = 12.2g/ 153.8g/mol = 0.07932mol

moles of Cl = 4 × 0.07932mol = 0.31728mol

1mole = 6.022 ×1023 atoms

number of Cl atoms = (6.022×1023atoms/1mol) × 0.31728mol

number of Cl atoms = 1.91×1023atoms

f)

Number of moles of HNO3 = 4.20g/63.0g/mol = 0.06667mol

Number of molecules of HNO3 = 6.022× 1023 molecles /mol × 0.06667mol

Number of molecules of HNO3 = 4.01× 1022molecules